# Polarization Signatures of Quasi-Periodic Oscillations in Simulated Tilted, Truncated Disks

### v2

@@ Abstract
We utilize the Monte-Carlo radiation transport code, Pandurata, to create images, spectra, polarization maps, and light curves from a set of general relativistic magnetohydrodynamic simulations of tilted, truncated, black hole accretion disks. Truncation can have spectral and polarization signatures all its own; tilt introduces both inclination and azimuthal dependencies into the spectra and polarization; and precession and oscillations of the tilted accretion flow inside the truncation radius (what we posit to be the “corona”) introduce time dependencies or periodicity to all of this. We use the ray-traced results from our simulations to evaluate the feasibility of measuring these effects, particularly in the context of current and future X-ray polarization observatories. Such detections could greatly improve our understanding of the geometry of accretion disks and coronae in the hard state, the physics of quasi-periodic oscillations (QPOs), and how system properties evolve as sources approach the hard-to-soft state transition.
Accretion (14) — X-ray polarization(1571) — Relativistic disks (1388) — Rotating black holes (1406) — Low-mass x-ray binary stars (939)
@@ 1 Introduction
There are two main contributions to the X-ray spectra of BHXRBs that control their hardness during an outburst – a soft, thermal component (below approximately 3 keV) and a hard, power-law component (extending roughly between 10 - 100 keV). The soft component is attributed to thermal, blackbody-like radiation coming from a relatively cold, optically thick and geometrically thin disk. The power-law component is then generated by the inverse-Compton scattering of soft seed photons from the disk by a radiatively inefficient cloud of hot electrons, i.e., the “corona.” The location and geometry of the corona are still under debate. We favor a truncated-disk geometry with a standard, thin disk truncated outside the last stable orbit, and the corona as a geometrically thick, radiatively inefficient accretion flow filling in the gap (Eardley & Lightman, 1975; Esin et al., 1997; Liu et al., 2007).
@@
A promising model for this QPO is having a finite, inner region of the accretion flow precess as a rigid body (Schnittman et al., 2006a; Ingram et al., 2009). This model can tie in with the truncated disk picture of the hard state if it is the corona that is precessing (Ingram & Done, 2011; Fragile et al., 2016). This naturally explains the rise in the QPO frequency during the early phases of the outburst cycle (when the truncation radius is moving inward; Ingram & Done, 2011; Fragile et al., 2016; Kubota et al., 2024) and easily reproduces the inclination dependence seen in QPO observations (Veledina et al., 2013). Thus, the time variability and geometry of black hole accretion flows may be closely linked, and polarization measurements may be able to shed light on both.
@@ 2.1 Truncated Disk Simulations
We consider three, fully 3D GRMHD simulations of accretion disks. These include the a9b0L4 and a9b15L4 simulations from Bollimpalli et al. (2024) and the isolated, tilted torus simulation from Appendix A of that same paper. The first two simulations differ in terms of the initial alignment of angular momentum axes between the disk and the black hole (with dimensionless spin parameter $a_{*}=a/M=0.9$ in both cases). In simulation a9b0L4, the axes are aligned, while in the a9b15L4 and isolated torus simulations, the initial disk angular momentum vectors are oriented along the $z$-axis, while the black hole spin axis is tilted by an angle $\beta_{0}=15^{\circ}$ toward the $-x$-axis.
All three simulations are initialized with a finite torus having an inner edge at $6.5\,r_{g}$ and pressure maximum at $9\,r_{g}$, where $r_{g}\equiv GM/c^{2}$. In the a9b0L4 and a9b15L4 simulations, this torus is surrounded by a thin slab of gas with a constant height of $H=0.4\,r_{g}$, extending from $15\,r_{g}$ to the outer boundary of the simulation domain at approximately $250\,r_{g}$. To preserve the desired thin structure of the outer disk, we implement an artificial cooling function, which is applied exclusively to radii beyond $15\,r_{g}$.
@@
[…]
@@
The results of these simulations are fully described in Bollimpalli et al. (2023, 2024). The most relevant findings for our current purposes are: 1) that there is a clear signature of solid-body precession of the inner torus when it is tilted, while there is no precession in the outer, thin disk (once the initial bending wave passes by a given radius); and 2) that the tilted simulations also exhibit high-frequency oscillations. Our goal in this paper is to determine whether signatures of tilted, truncated disks and QPOs might be observable in polarized light by IXPE or XPoSat.
@@ 2.2 Pandurata
We use the Monte-Carlo radiation transport code Pandurata  (Schnittman & Krolik, 2013) to post-process our numerical simulations and create images, lightcurves, spectra, and polarization maps analogous to X-ray polarization observations.
@@
The first step in analyzing the simulation data is to convert from “code” to cgs units. We do so by specifying a black hole mass ($10M_{\odot}$ in this case) and Eddington-scaled accretion rate (chosen to be $\dot{M}\approx 0.1L_{\mathrm{Edd}}/c^{2}$), as in Schnittman & Krolik (2013). After reading in and remapping the simulation data, the next step is to locate the photosphere of the disk. We do this by integrating the quantity $\kappa\rho$ starting from the pole of the grid and proceeding toward the midplane following a constant coordinate radius $r$ and azimuth $\phi$ until $\tau=2$, where
[…]
and $\kappa=0.4$ cm2 g-1 is the scattering opacity. The functions $\tau_{<}(\theta)$ and $\tau_{>}(\theta)$ can be thought of as the optical depth to escape from the bottom or top of the disk, respectively. Figure 1 shows the photosphere for one azimuth near the end of our untilted and tilted simulations. Interestingly, even though we are treating the hot, thick, inner torus region as a proxy for the corona, it is dense enough in our simulations to be optically thick (exist beneath the photosphere).
@@
All the Cosmos++ simulations use a $\Gamma=5/3$ polytrope equation of state, corresponding to a non-relativistic ideal gas-like fluid. This is most appropriate for the coronal regions, where we can directly relate the code temperature to the electron temperature. For the optically thick disk, we instead interpret the code pressure $P$ as representing a radiation-pressure-dominated fluid with
$T_{\rm rad}=\left(\frac{3c}{4\sigma}P\right)^{1/4}\,,$ (3)
and then using this temperature for the thermal seed spectrum $T_{\rm eff}=T_{\rm rad}$. This same procedure was used in the very first demonstration of the ray tracing of a GRMHD simulation (Schnittman et al., 2006b).
Starting from the photosphere at each radius and azimuth, Pandurata launches $\sim 10^{4}$ packets of thermal photons into the optically thin regions surrounding the disk [^1] To test convergence, we redid a couple of our Pandurata runs using four times as many photon packets. Those results agreed to within 10% or less of the original.. The initial direction of a photon packet is selected from an isotropic distribution in the emitting fluid frame, limited by the photosphere surface. Each packet is weighted by a number of geometric emission factors, such that a photon packet emitted from a small patch of optically thick, scattering-dominated accretion disk would have a spectrum of
@@
where $F_{\nu}$ has units of spectral luminosity [erg s-1 Hz-1], $f$ is a hardening factor (we use 1.8), $B_{\nu}$ is the usual blackbody function, $T_{\mathrm{eff}}$ is the effective local temperature, $\cos\theta_{\mathrm{em}}$ is a geometric factor, $f_{\mathrm{limb}}$ treats the limb darkening (Chandrasekhar, 1960), $dA$ is the proper area of the emission region, $d\Omega$ is the solid angle, and $1/u^{t}=d\tau/dt$ converts from time in the emission frame to that of the coordinate frame. For $T_{\mathrm{eff}}(r,\phi)$, we take an average of the radiation temperature from the top and bottom of the photosphere, assuming the gas and radiation are in pressure equilibrium.
The photon packets then follow null geodesics, using the same integrator as described in Schnittman & Bertschinger (2004). Because the geodesic trajectories are independent of photon energy, we simultaneously track photons covering energies in the range $[10^{-3}-10^{4}]$ keV. However, we do still need to account for Doppler and gravitational redshifting. To do so, we track each photon packet’s frequency (or energy) in both the “emitter” and “observer” frames. For a photon packet emitted in a frame with fluid four-velocity $u^{\mu}$(em) and photon four-momentum $k_{\mu}$(em) and observed in a frame with $u^{\nu}$(obs) and $k_{\nu}$(obs), the redshifted frequency is
@@
Whenever a photon packet scatters off the disk or an electron in the corona, the frequencies $\nu_{i}$ are updated and the old observed frame becomes the new emitted frame.
@@
To track the polarization along the geodesic path, we take advantage of the fact that the polarization vector $\mathbf{f}$ is a space-like vector normal to the photon wave vector $\mathbf{k}$. Therefore, it is constrained by $\mathbf{f}\cdot\mathbf{f}=1$ and $\mathbf{f}\cdot\mathbf{k}=0$ (Connors et al., 1980). Since $\mathbf{k}$ is a null vector, we can always redefine $\mathbf{f}$ by adding some multiple of $\mathbf{k}$ to it and thus write the polarization vector as
@@
Once the photon packets have finally escaped the simulation domain on their way to an observer, they are collected into $\{\theta,\phi\}$ bins. We use 41 evenly spaced bins in the $\theta$-direction and 8 in the $\phi$-direction [^2] For some of the untilted simulation results, we collect all of the photon packets into a single azimuthal bin, as appropriate in the case of axisymmetry.. When a photon packet reaches “infinity”, or in our case a spherical shell with $R_{\rm shell}=10000\,r_{g}$, the detector coordinates are defined such that the $+x$ axis is along the $+\phi$ direction and the $+y$ axis is in the $-\theta$ direction. This corresponds to horizontal polarization having $\psi=0^{\circ}$ or $180^{\circ}$, and vertical polarization characterized by $\psi=\pm 90^{\circ}$. Furthermore, for these calculations, we employ the “fast light” approximation which does not account for light travel time across the simulation volume. This is appropriate for simulations where the data is stored and analyzed with relatively coarse temporal sampling.
@@ 3.1 Untilted, Truncated Disk
Since the azimuthal angle is not a factor in this section, we focus instead on the inclination-dependence of our results. We adopt the usual convention, with $i=0^{\circ}$ being a view straight down the black hole spin axis (face-on to the disk) and $i=90^{\circ}$ being an observer in the plane of the disk (edge-on). Comparing a nearly face-on view with an edge-on one, we would expect the face-on one to have a higher overall flux (since it experiences less limb darkening), yet a lower amount of polarization (since photon packets leaving normal to the disk surface start with zero polarization and have a shorter path through the corona, thus reducing their likelihood of scattering). This is confirmed in Figure 2, where we plot flux and polarization spectra for observer inclinations of $i=18$, 45, and $72^{\circ}$. The nearly face-on disk ($i=18^{\circ}$) is approximately 10 times brighter across much of the spectrum than the nearly edge-on case ($i=72^{\circ}$), but only has a polarization degree of $<2$%, whereas the nearly edge-on case has a polarization of $\geq 5$%, except at high energies. In all cases, the polarization angle is close to $0^{\circ}$ for $E\lesssim 10$ keV, as expected for emission from a planar disk with its symmetry axis pointed vertically (Schnittman & Krolik, 2009).
@@
By comparing the total intensity and polarization images of the nearly face-on and edge-on cases as depicted in Figure 3, we can easily understand the results of Figure 2. For the nearly edge-on image (bottom panel), there is a high degree of scattering-induced limb darkening of the original photon packets (Chandrasekhar, 1960). Additionally, most of the photon packets must traverse a lengthy path through the corona, so are more likely to be scattered out of the observer’s line of sight or otherwise attenuated. By contrast, for the nearly face-on view (top panel), most of the photon packets start with relatively little limb darkening and travel a shorter path through the corona to the observer. This explains why the top image has a higher overall intensity. However, since photons launched perpendicular to the disk are created with very low polarization fractions, the top (nearly face-on) image shows mostly short polarization vectors, except close to the black hole where light-bending effects amplify the polarization. The nearly edge-on image, by contrast, shows much longer, mostly horizontal, polarization vectors, consistent with the fact that most of these photon packets were launched nearly parallel to the disk surface and had a higher likelihood of scattering within the disk (Chandrasekhar, 1960).
Figure 3: Total intensity images with polarization vectors overlaid for low ($i=18^{\circ}$, top panel) and high ($i=72^{\circ}$, bottom panel) inclinations. Both images represent $40r_{g}\times 40r_{g}$ fields of view centered on the black hole.

### v3

@@ Abstract
We utilize the Monte Carlo radiation transport code, Pandurata, to create images, spectra, polarization maps, and light curves from a set of general relativistic magnetohydrodynamic simulations of tilted, truncated, black hole accretion disks. Truncation can have spectral and polarization signatures all its own; tilt introduces both inclination and azimuthal dependencies into the spectra and polarization; and precession and oscillations of the tilted accretion flow inside the truncation radius introduce time dependencies or periodicity to all of this. We use the ray-traced results from our simulations to evaluate the feasibility of measuring these effects, particularly in the context of current and future X-ray polarization observatories. Such detections could greatly improve our understanding of the geometry of accretion disks and coronae in the hard state, the physics of quasi-periodic oscillations (QPOs), and how system properties evolve as sources approach the hard-to-soft state transition.
Accretion (14) — Starlight polarization (1571) — Relativistic disks (1388) — Rotating black holes (1406) — Low-mass x-ray binary stars (939)
@@ 1 Introduction
There are two main contributions to the X-ray spectra of BHXRBs that control their hardness during an outburst – a soft, thermal component (below approximately 3 keV) and a hard, power-law component (extending roughly between 10 - 100 keV). The soft component is attributed to thermal, blackbody-like radiation coming from a relatively cold, optically thick and geometrically thin disk. The power-law component is then generated by the inverse-Compton scattering of soft seed photons from the disk by a radiatively inefficient cloud of hot electrons called the “corona.” The location and geometry of the corona are still under debate. We favor a truncated-disk geometry with a standard, thin disk truncated outside the last stable orbit, and the corona as a geometrically thick, radiatively inefficient accretion flow filling in the gap (Eardley & Lightman, 1975; Esin et al., 1997; Liu et al., 2007).
@@
A promising model for this QPO is having a finite, inner region of the accretion flow precess as a rigid body (Schnittman et al., 2006a; Ingram et al., 2009). This model can tie in with the truncated disk picture of the hard state if it is the corona that is precessing (Ingram & Done, 2011; Fragile et al., 2016). This naturally explains the rise in the QPO frequency during the early phases of the outburst cycle when the truncation radius is moving inward (Ingram & Done, 2011; Fragile et al., 2016; Kubota et al., 2024) and easily reproduces the inclination dependence seen in QPO observations (Veledina et al., 2013). Thus, the time variability and geometry of black hole accretion flows may be closely linked, and polarization measurements may be able to shed light on both.
@@ 2.1 Truncated Disk Simulations
We consider three, fully 3D GRMHD simulations of accretion disks. These include the a9b0L4 and a9b15L4 simulations from Bollimpalli et al. (2024) and the isolated, tilted torus simulation from Appendix A of that same paper. The first two simulations differ in terms of the initial alignment of angular momentum axes between the disk and the black hole (with dimensionless spin parameter $a_{*}=a/M=0.9$ in both cases). In simulation a9b0L4, the axes are aligned, while in the a9b15L4 and isolated torus simulations, the initial disk angular momentum vectors are oriented along the $z$-axis, while the black hole spin axis is tilted by an angle $\beta_{0}=15^{\circ}$ toward the x-x-axis.
All three simulations are initialized with a finite torus having an inner edge at $6.5\,r_{g}$ and pressure maximum at $9\,r_{g}$, where $r_{g}\equiv GM/c^{2}$. In the a9b0L4 and a9b15L4 simulations, this torus is surrounded by a thin slab of gas with an initial height of $H=0.4\,r_{g}$, extending from $15\,r_{g}$ to the outer boundary of the simulation domain at approximately $250\,r_{g}$. To preserve the desired thin structure of the outer disk, we implement an artificial cooling function, which is applied exclusively to radii beyond $15\,r_{g}$.
@@
The simulations are performed in modified Kerr-Schild, spherical-polar coordinates, enabling the placement of the inner domain boundary inside the black hole event horizon. The grid is logarithmically spaced in $r$, covering the range $r\in[1.4\,r_{g},\,40^{1.5}\,r_{g}]$, and uniformly spaced in $\phi$, encompassing the entire $0,2\pi]$ domain. The polar angle $\theta$ is related to the uniformly-spaced coordinate $x_{2}$, with the grid spacing defined as
@@
The results of these simulations are fully described in Bollimpalli et al. (2023, 2024). The most relevant findings for our current purposes are: 1) that there is a clear signature of solid-body precession of the inner torus when it is tilted, while there is no precession in the outer, thin disk (once an initial bending wave passes by a given radius); and 2) that the tilted simulations also exhibit high-frequency oscillations. Our goal in this paper is to determine whether signatures of tilted, truncated disks and QPOs might be observable in polarized light by IXPE or XPoSat.
@@ 2.2 Pandurata
We use the Monte Carlo radiation transport code Pandurata  (Schnittman & Krolik, 2013) to post-process our numerical simulations and create images, lightcurves, spectra, and polarization maps analogous to X-ray polarization observations.
@@
The first step in analyzing the simulation data is to convert from “code” to cgs units. We do so by specifying a black hole mass ($10M_{\odot}$ in this case) and a density sufficient to produce a luminosity of a few percent of Eddington, as in Schnittman & Krolik (2013). After reading in and remapping the simulation data, the next step is to locate the photosphere of the disk. We do this by integrating the quantity $\kappa\rho$ starting from the pole of the grid and proceeding toward the midplane following a constant coordinate radius $r$ and azimuth $\phi$ until $\tau=2$, where
$\tau_{<}(\theta)$ $=$ $-\int^{\theta}_{\pi}u^{t}\kappa\rho\sqrt{g_{\theta\theta}}d\theta~,$
$\tau_{>}(\theta)$ $=$ $\int^{\theta}_{0}u^{t}\kappa\rho\sqrt{g_{\theta\theta}}d\theta~,$ (2)
and $\kappa=0.4$ cm2 g-1 is the scattering opacity. The functions $\tau_{<}(\theta)$ and $\tau_{>}(\theta)$ can be thought of as the optical depth to escape from the bottom or top of the disk, respectively. Figure 1 shows the photosphere for one azimuth near the end of our untilted and tilted simulations. We point out that our inner torus is dense enough to be optically thick (exist beneath the photosphere), so although we have a truncated disk in the sense that we have two hydrodynamically distinct regions to our disk, our inner torus is not a good proxy for the corona. Instead, the corona is found above and below the disk, as indicated by the colored regions in Figure 1.
@@
The particular Cosmos++ simulations being studied here do not include radiative transport, instead using a $\Gamma=5/3$ polytropic equation of state. The only cooling that is allowed (aside from adiabatic expansion) is the artificial cooling function used to keep the outer disk thin. As a result, the background gas, which can also be affected by numerical floors, can reach unrealistically high temperatures ($10^{9}$ K). Since the temperature of the background gas (really the electrons) strongly influences the inverse Compton scattering in Pandurata, we need to rescale the background temperatures. We apply a power-law scaling that effectively caps the background temperatures at $10^{8}$ K. For the photosphere temperature, we recover it by interpreting the code pressure $P$ as representing a radiation-pressure-dominated fluid with
$T_{\rm rad}=\left(\frac{3c}{4\sigma}P\right)^{1/4}\,.$ (3)
This temperature is then used for the thermal seed spectrum $T_{\rm eff}=T_{\rm rad}$. This same procedure was used in the very first demonstration of the ray tracing of a GRMHD simulation (Schnittman et al., 2006b).
Starting from the photosphere at each radius and azimuth, Pandurata launches $\sim 10^{4}$ packets of thermal photons into the optically thin regions surrounding the disk [^1] To test convergence, we redid a couple of our Pandurata runs using four times as many photon packets. Those results agreed to within 10% or less of the original.. The initial direction of a photon packet is selected from an isotropic distribution in the emitting fluid frame, limited by the photosphere surface. Each packet is weighted by a number of geometric emission factors, such that a photon packet emitted from a small patch of optically thick, scattering-dominated accretion disk should have a spectrum in the local fluid frame of
@@
where $F_{\nu}$ has units of spectral luminosity [erg s-1 Hz-1], $f$ is a hardening factor (we use 1.8), $B_{\nu}$ is the usual blackbody function, $T_{\mathrm{eff}}$ is the effective local temperature, $\theta_{\mathrm{em}}$ is the emission angle relative to the disk normal, $f_{\mathrm{limb}}$ treats the limb darkening (Chandrasekhar, 1960), $dA$ is the proper area of the emission region, $d\Omega$ is the solid angle, and $1/u^{t}=d\tau/dt$ converts from time in the emission frame to that of the coordinate frame. For $T_{\mathrm{eff}}(r,\phi)$, we take an average of the radiation temperature from the top and bottom of the photosphere, assuming the gas and radiation are in pressure equilibrium.
The photon packets then follow null geodesics, using the same integrator as described in Schnittman & Bertschinger (2004). Because the geodesic trajectories are independent of photon energy, we simultaneously track photons covering energies in the range $10^{-3}-10^{4}]$ keV. However, we do still need to account for Doppler and gravitational redshifting. To do so, we track each photon packet’s frequency (or energy) in both the “emitter” and “observer” frames. For a photon packet emitted in a frame with fluid four-velocity $u^{\mu}$(em) and photon four-momentum $k_{\mu}$(em) and observed in a frame with $u^{\nu}$(obs) and $k_{\nu}$(obs), the redshifted frequency is
@@
Whenever a photon packet scatters off the disk or an electron in the corona, the frequencies $\nu_{i}$ are updated and the old observer frame becomes the new emitter frame.
One of the most notable features of the Pandurata  approach to radiation transport is the treatment of photon packets, which allow us to treat a broad spectrum of photons with a single geodesic trajectory. When including electron scattering, we are forced to make an approximation that all photons scatter with identical cross section, i.e. the Thomson scattering distribution. At very high energies ($\gtrsim 100$ keV), it would be more appropriate to apply a Klein-Nishina cross section, but this would require energy-dependent scattering, a feature that only slows down the ray-tracing calculation. Rather, we choose to focus on Thomson scattering, with a thermal Maxwell-Jutner distribution for the coronal electrons, with the temperature taken from the GRMHD simulations as described above. Since we do not report results above 100 keV, this should be a reasonable approach.
After a scattering event, the entire spectrum of the photon packet shifts up or down, according to the classical (inverse) Compton scattering rules. We are particularly careful to modify the spectrum (in units of erg/s/cm2/Sr/Hz) so that photon number is conserved, as expected for scattering events. When scattering off the thermal accretion disk, we treat it as a semi-infinite scattering atmosphere and take the median number of events as 2 (Chandrasekhar, 1960), adjusting the spectrum accordingly.
@@
To track the polarization along the geodesic path, we take advantage of the fact that the polarization 4-vector $\mathbf{f}$ is a space-like vector normal to the photon wave vector $\mathbf{k}$. Therefore, it is constrained by the normalizations $\mathbf{f}\cdot\mathbf{f}=1$ and $\mathbf{f}\cdot\mathbf{k}=0$ (Connors et al., 1980). Since $\mathbf{k}$ is a null vector, we can always redefine $\mathbf{f}$ by adding some multiple of $\mathbf{k}$ to it, thus writing the polarization vector as
@@
As described in Schnittman & Krolik (2013), the polarization vector can be determined at any point along a geodesic path in the Kerr geometry using Boyer-Lindquist coordinates via the complex-valued Walker-Penrose constant $\kappa_{\rm wp}$ (Walker & Penrose, 1970; Connors & Stark, 1977). Given the wave vector $k^{\mu}$ along the geodesic path, $\kappa_{\rm wp}$ is computed at any point by
$\kappa_{\rm wp}$ $=$ $\left\{(k^{t}f^{r}-k^{r}f^{t})+a\sin^{2}\theta(k^{r}f^{\phi}-k^{\phi}f^{r})\right$. (8)
$-i[(r^{2}+a^{2})(k^{\phi}f^{\theta}-f^{\phi}k^{\theta})$
$\left.-a(k^{t}f^{\theta}-k^{\theta}f^{t})]\sin\theta\right\}(r-ia\cos\theta)\,.$
Combined with the above normalizations $\mathbf{f}\cdot\mathbf{f}=1$ and $\mathbf{f}\cdot\mathbf{k}=0$, we have four linear equations for the four components of $f^{\mu}$.
Once the photon packets have finally escaped the simulation domain on their way to an observer, they are collected into $\{\theta,\phi\}$ bins. We use 41 bins evenly spaced in $\cos\theta$ and 8 evenly spaced in $\phi$ [^2] For some of the untilted simulation results, we collect all of the photon packets into a single azimuthal bin, as appropriate in the case of axisymmetry.. When a photon packet reaches “infinity”, or in our case a spherical shell with $R_{\rm shell}=10,000\,r_{g}$, the detector coordinates are defined such that the +x+x axis is along the $\phi$ direction and the +y+y axis is in the $\theta$ direction. This corresponds to horizontal polarization having a polarization angle $\psi=0^{\circ}$ or $180^{\circ}$, and vertical polarization characterized by $\psi=\pm 90^{\circ}$. Furthermore, for these calculations, we employ the “fast light” approximation which does not account for light travel time across the simulation volume. This is appropriate for simulations where the data is stored and analyzed with relatively coarse temporal sampling.
@@ 3.1 Untilted, Truncated Disk
Since the azimuthal angle is not a factor in this section, we focus instead on the inclination-dependence of our results. We adopt the usual convention, with $i=0^{\circ}$ being a view straight down the black hole spin axis (face-on to the disk) and $i=90^{\circ}$ being an observer in the plane of the disk (edge-on). Comparing a nearly face-on view with a nearly edge-on one, we would expect the face-on one to have a higher overall flux (since it experiences less limb darkening), yet a lower amount of polarization (since photon packets leaving normal to the disk surface start with zero polarization and have a shorter path through the corona, thus reducing their likelihood of scattering). This is confirmed in Figure 2, where we plot flux and polarization spectra for observer inclinations of $i=18$, 45, and $72^{\circ}$. The nearly face-on disk ($i=18^{\circ}$) is approximately 5 times brighter across much of the spectrum than the nearly edge-on case ($i=72^{\circ}$), but only has a polarization degree of <2<2% (except at high energies), whereas the nearly edge-on case has a polarization of $\geq 2$%. In all cases, the polarization angle is close to $0^{\circ}$ (or $180^{\circ}$) for $E\lesssim 1$ keV, as expected for emission from a planar disk with its symmetry axis pointed vertically (Schnittman & Krolik, 2009) [^3] The discontinuous jumps in polarization angle in spectra like Fig. 2 come from the fact that angles of $0^{\circ}$ and $180^{\circ}$ are degenerate with each other, so values can quickly swing from one limit to the other. Other sharp jumps occur at high energies where our photon statistics are poor..
@@
By comparing the total intensity and polarization images of the nearly face-on and edge-on cases as depicted in Figure 3, we can easily understand the results of Figure 2. For the nearly edge-on image (bottom panel), there is a high degree of scattering-induced limb darkening of the original photon packets (Chandrasekhar, 1960). Additionally, most of the photon packets must traverse a lengthy path through the corona, so are more likely to be scattered out of the observer’s line of sight or otherwise attenuated. By contrast, for the nearly face-on view (top panel), most of the photon packets start with relatively little limb darkening and travel a shorter path through the corona to the observer. This explains why the top image has a higher overall intensity. However, since photons launched perpendicular to the disk are created with very low polarization fractions, the top (nearly face-on) image shows only moderate length polarization vectors. The nearly edge-on image, by contrast, shows much longer polarization vectors, consistent with the fact that most of these photon packets had a higher likelihood of scattering within the corona and picking up polarization.
Figure 3: Total intensity images with polarization vectors overlaid for low ($i=18^{\circ}$, top panel) and high ($i=72^{\circ}$, bottom panel) inclinations for our untilted, truncated disk simulation (a9b0L4). Both images represent $40r_{g}\times 40r_{g}$ fields of view centered on the black hole.